OpenStudy (anonymous):

I'm deperately trying to solve find the vertical asymptotes: (x^3+2x^2-20x+24)/(x^2-x-12)

OpenStudy (unklerhaukus):

factorise the numerator and the denominator

OpenStudy (lgbasallote):

the numerator?

OpenStudy (anonymous):

Set denominator =0.

OpenStudy (anonymous):

I tried, I came up with x + 3 - [ (5x+60)/(x^2-x-12) ]

OpenStudy (unklerhaukus):

\[\frac{x^3+2x^2-20x+24}{x^2-x-12}=\frac{(\quad)(\quad)(\quad)}{(\quad)(\quad)}\]

OpenStudy (anonymous):

(Yes, must factor numerator too to see if any factors cancel out, leaving only holes and not asymptotes)

OpenStudy (anonymous):

Yes, I have holes. Is it possible that there are simply no V.A.s?

OpenStudy (anonymous):

Thanks for the help so far

OpenStudy (unklerhaukus):

\[x^2-x-12=(x+\quad)(x-\quad)\]

OpenStudy (anonymous):

(x-4)(x+3)

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

ans when I solve, they both are = 0. So, therefore, no V.A. right?

OpenStudy (unklerhaukus):

the product of two terms is zero when EITHER is zero

OpenStudy (anonymous):

Right, so when I plug the numbers back into the denominator I get 0. I'm basically just trying to see if it's possible to have no vertical asymptote

OpenStudy (anonymous):

because I got that x+3 is the slant asymptote. I hope that's correct

OpenStudy (unklerhaukus):

but we should try to factorise the numerator to see if any factors cancel out, like CliffSedge said

OpenStudy (unklerhaukus):

\[x^3+2x^2-20x+24=(x+\quad)(x-\quad)(x-\quad)\]

OpenStudy (anonymous):

\[x+3 - \frac{ 5x+60 }{ x^2 - x - 12 }\] this is what I got using using long division

ganeshie8 (ganeshie8):

then your slant asymptote would be y = x+3, but still vertical asymptotes exist, slant asymptotes dont disturb VA i think, Unklw may correct me

OpenStudy (anonymous):

Yeah, thanks for your input. The vertical asymptotes should be able to be found by factoring the denominator of the function. So it doesn't disturb it. I'm just wondering why both factors equal zero when I plug them into the denominator of the function

ganeshie8 (ganeshie8):

hmm you asking about "zero product prperty" ?

OpenStudy (anonymous):

If it is equal to zero, then it is not a vertical asymptote. So i'm wondering how it's possible to have no V.A.

ganeshie8 (ganeshie8):

its the opposite of what u think actually :) precisely when something in the denominator equals 0, we get a VA or hole.

ganeshie8 (ganeshie8):

1/0.00000000000000000000001 = Very BIG

ganeshie8 (ganeshie8):

when denominator approaches 0, the value of function becomes BIGG

ganeshie8 (ganeshie8):

as it approaches 0, the function increases without any limit... thats where we get VA

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

so (x-4), VA = 4, (x+3), VA = -3

ganeshie8 (ganeshie8):

vertical asymptote is a "line"

OpenStudy (anonymous):

a slanted asympoted right?

OpenStudy (anonymous):

x+3

ganeshie8 (ganeshie8):

so, we say, (x-4) = 0 : x=4 is VA or hole (x+3) = 0 : x = -3 is VA or hole

ganeshie8 (ganeshie8):

a slanted asymtopte is also a line

ganeshie8 (ganeshie8):

we got the quotient (x+3) right ? so we say, the rational function has slanted asymtopte : y = x+3

OpenStudy (anonymous):

Okay, I see. Thank you!

ganeshie8 (ganeshie8):

again, we didnt knw whether our denum factors cancel out, as @UnkleRhaukus was saying, you need to factor numerator and see if any common factors exist... if so, then we wil get holes

OpenStudy (anonymous):

the only factoring I could do is -5(x+12)

ganeshie8 (ganeshie8):

yeah so no holes. the two discontinuties which we have found, x=4 and x=-3 are VAs !

OpenStudy (anonymous):

Well, thank you :) I'd love help for the rest of this problem..if ur up for it. :p

OpenStudy (anonymous):

I have to find x-ints, y-ints, etc.

ganeshie8 (ganeshie8):

those are easy ones

ganeshie8 (ganeshie8):

to find y-int, put x=0 in the rational function

OpenStudy (anonymous):

found Y to be -2. X seems a little harder...

ganeshie8 (ganeshie8):

f(x) = \(\frac{x^3+2x^2-20x+24}{x^2-x-12} \) f(0) = ?

ganeshie8 (ganeshie8):

ok good :) finding X is bit tricky

ganeshie8 (ganeshie8):

for finding x-int, you set numerator = 0

ganeshie8 (ganeshie8):

x^3+2x^2-20x+24 = 0

ganeshie8 (ganeshie8):

we need to factor !

ganeshie8 (ganeshie8):

(x - ) (x - ) (x - ) = 0

OpenStudy (anonymous):

been trying for a couple of minutes now lol

ganeshie8 (ganeshie8):

you familiar wid rational roots theorem ?

OpenStudy (anonymous):

Sounds familair

OpenStudy (anonymous):

I'm going to look it up

ganeshie8 (ganeshie8):

ok, but il explain u briefly quickly

ganeshie8 (ganeshie8):

we want to find roots of this : x^3+2x^2-20x+24 = 0 rational root theorem says, the rational roots of a polynomial will always come from : 1) factors of 24 : \(\pm1 , \pm2, \pm3,....\) 2) factors of 1: \(\pm1\)

ganeshie8 (ganeshie8):

rational roots : p/q, we pick p from 1) q from 2)

ganeshie8 (ganeshie8):

see if that makes sense, see ur notes also hmm

OpenStudy (anonymous):

Ahh, yeah I do recognize that.